# BERTs are Generative In-Context Learners

### v1

@@ Abstract
This paper explores the in-context learning capabilities of masked language models, challenging the common view that this ability does not ‘emerge’ in them. We present an embarrassingly simple inference technique that enables DeBERTa to operate as a generative model without any additional training. Our findings demonstrate that DeBERTa can match and even surpass GPT-3, its contemporary that famously introduced the paradigm of in-context learning. The comparative analysis reveals that the masked and causal language models behave very differently, as they clearly outperform each other on different categories of tasks. This suggests that there is great potential for a hybrid training approach that takes advantage of the strengths of both training objectives.
@@ 1 Introduction
Masked language models used to dominate the field of natural language processing due to their adaptability across diverse tasks and their superior performance compared to causal language models (Radford et al., 2018; Devlin et al., 2019). Between 2018 and 2020, the field witnessed a surge in the development of these models (Devlin et al., 2019; Liu et al., 2019; Lan et al., 2020, inter alia). However, the introduction of GPT-3 in Brown et al. (2020) marked a paradigm shift with its demonstration of in-context learning, where a model infers a task from context without any finetuning. Thus, avoiding the need for extensive hand-annotated datasets and the need for any deep-learning expertise, it is particularly attractive for practical applications, resulting in a drastic shift away from BERT-like models.
@@
Figure 1: The average 1-shot performance across four groups of NLP tasks. We compare the scaling abilities of DeBERTa (four sizes in red) with GPT-3 (eight sizes in blue). Even though these models rely on different training objectives, they scale in a similar log-linear manner overall. Yet, on a task-by-task basis, the pretraining methods lead to substantial differences between them.
@@
Previous studies of ‘emergent’ in-context-learning abilities often focus only on causal language models (Saunshi et al., 2021; Olsson et al., 2022; Wei et al., 2022; Wang et al., 2023, inter alia). However, this paper demonstrates that in-context learning is a more general phenomenon and should not be studied with a singular pretraining objective in mind. Moreover, the assumed inability of masked language models to perform (generative) in-context learning has rendered them outdated – as explicitly noted by Tay et al. (2023): “BERT-style models are very restricted in their generative capabilities. Because of the cumbersomeness of task-specific classification heads, we strongly do not recommend using this class of autoencoding models moving forward and consider them somewhat deprecated.”
@@
In this paper, we challenge these prevailing assumptions about masked language models (MLMs). We present empirical evidence showing that DeBERTa, an MLM released just one month after GPT-3, is equally adept at in-context learning. Our findings suggest that the capacity for in-context learning is not tied to the training objective, but can be achieved across different types of language models. Interestingly, we show that DeBERTa does not simply mimic the performance of GPT-3 – it is clearly much better on tasks such as language understanding, and, on the other hand, much worse on tasks such as closed-book question answering. This suggests that masked and causal language modeling are two complementary training objectives and that there is a great potential for a training method that combines the strengths of both objectives. Finally, scaling (performance improvement with increased size of pretrained language models) is a crucial feature of modern language models, we demonstrate that MLMs do scale on in-context learning (Figure 1).
@@
We introduce a simple inference technique that transforms an MLM into a generative model without any further training. Using publicly available DeBERTa checkpoints, we show that the MLM training objective not only provides a versatile way of encoding text, but is also competitive in text generation and text completion ranking. This claim is tested by following the same evaluation suite as GPT-3, speculating on an ‘alternative reality’ in which a masked language model is the first model reported to achieve the so-called ‘emergent’ in-context learning abilities.
@@
Figure 2: Illustration of the proposed methods for using a masked language model for text generation and text ranking, by reformatting the sequence of input tokens and without any additional training. The left figure shows how to modify the input tokens to autoregressively generate text. The right figure illustrates a similar approach that is used to retrieve a pseudo-log-likelihood score of a text sequence that can be used to rank multiple sequences by their joint probabilities.
@@ 2.1 Text generation
This straightforward inference scheme would be enough if the pretraining process was designed with this use case in mind. However, since our goal is to repurpose an existing masked language model, we have to complicate the method with two modifications that are also illustrated in Figure 2:
@@
    However, the addition of this end-of-sequence token creates a problem – it raises the probability that the masked token should end the sequence (for example with a full stop). Thus, in order to obtain a less restricted continuation, we include additional [MASK] tokens to pad the space in front of the end-of-sequence token. Specifically, we use two additional masks for the DeBERTa models. [^1] Note that this is not an arbitrary number but it is model-specific – DeBERTa models were pretrained to unmask spans of masked tokens where the longest allowed spans are three tokens long (He et al., 2021).
@@
While this method works with the same quadratic (in sequence length) time and memory complexity, it is slower in practice because it is not possible to cache the intermediate self-attention key and value vectors. Instead, the hidden representations of the whole sequence have to be recomputed in every step due to the bidirectional nature of the model. We believe that this limitation can be fixed, either by using prefix language modeling or by not updating hidden vectors of distant tokens every step. We leave these optimizations for future work.
@@ 2.2 Ranking
While this equation matches the training objective of causal language models, it is not suitable for masked language models because they are not trained to estimate $\operatorname{\mathbb{P}}(w_{i}\,|\,c\oplus w_{0}\dots w_{i-1})$. Instead, Wang and Cho (2019) proposed to modify Equation 1 to make it more appropriate for BERT-like models. Salazar et al. (2020) then empirically showed that the resulting pseudo-log-likelihood (PLL) score can be used to accurately rank text sequences by their likelihood. More specifically, the PLL score is approximately proportional to the conditional probability of a text sequence and is computed as follows:
@@
However, this approximation gets very inaccurate when there are strong local dependencies between tokens. As a counterexample, the estimated likelihood of the multi-token word ‘supercalifragilisticexpialidocious’ is seven orders of magnitude higher than that of the single-token word ‘super’, which is clearly an incorrect estimation of the relative frequencies of these words. [^2] This is because the tokenizer of DeBERTa splits the long word into 9 subwords and each of them is assigned an almost certain likelihood given the bidirectional context of surrounding subwords. The largest 1.4B DeBERTa estimates the pseudo log-likelihood of the first word to $-2.1$ while the second (very common) word has pseudo log-likelihood of $-9.6$.
@@
Figure 3: Length generalization measured with a ‘needle in a haystack’ benchmark. The $x$-axis indicates the total size of the ‘haystack’ and the $y$-axis indicates the position of the ‘needle’; the values show the average exact-match accuracy for a particular configuration. Unfortunately, GPT-3 is a closed-source model and the original version is not accessible, so we use an open-source replication of GPT-3, OPT by Zhang et al. (2022), which should perform similarly on this task.
@@ 2.3 Length generalization
A potentially limiting factor of using BERT-like models is that they are typically pretrained on shorter sequences than causal language models – arguably because the training of modern causal models is already optimized for in-context learning, which requires processing of long few-shot prompts. DeBERTa is not an exception of such pretraining; it was only trained with a relatively short maximum sequence length of 512 tokens (He et al., 2021). Fortunately, the architecture of DeBERTa can easily process much longer sequences than seen during training due to its use of relative positional embeddings with logarithmic buckets (Raffel et al., 2020).
We measure the extent to which DeBERTa generalizes to longer sequences with the ‘needle in a haystack’ test from RULER (Hsieh et al., 2024). Specifically, in our formulation of this task, a random 6-digit number (needle) is hidden in a long collection of essays (haystack). We then measure the exact-match accuracy of retrieving the hidden number given two variables: the total sequence length and the position of the needle in the haystack (more details about the evaluation setup are given in Section A.1). The results in Figure 3 demonstrate that DeBERTa processes sequences within GPT-3’s limit reasonably well and that it can generalize even beyond that length. In practice, this observation means that DeBERTa should be able to handle as many task demonstrations as GPT-3.
@@ 3 DeBERTa family of language models
This study uses the largest openly available English masked language model, DeBERTa (1.4B), and its smaller configurations (He et al., 2021). DeBERTa is an improved version of a BERT language model (Devlin et al., 2019) that uses an advanced attention mechanism with relative positional embeddings – apart from being trained on a larger corpus with a larger number of training steps.
@@
Compared to GPT-3 and modern large language models, DeBERTa was pretrained on a relatively small and clean text corpus – totalling 78GB of data after deduplication, the corpus is comprised of the English Wikipedia (12GB), BookCorpus (6GB; Zhu et al., 2015), OpenWebText (38GB; Gokaslan and Cohen, 2019), and STORIES (31GB; Trinh and Le, 2019). This is almost an order of magnitude less data than what was used to pretrain GPT-3, which uses 570GB of filtered CommonCrawl, WebText2 (roughly 26GB), two web-scraped book corpora (roughly 17GB and 76GB), and the English Wikipedia (roughly 4GB, estimated from Brown et al. (2020)).
@@
Interestingly, even though DeBERTa uses a substantially smaller training corpus, it is trained on more than three times more tokens than GPT-3 (1 trillion compared to 300 billion). However, the loss is computed only on 15% of tokens (150 billion) and it is not clear what would be the effective number of tokens used for pretraining. Nevertheless, the total compute used for training depends on the number of input tokens and it is roughly $8.0\cdot 10^{21}$ FLOPs for the 1.4B DeBERTa and $2.4\cdot 10^{21}$ FLOPs for the 1.3B GPT-3.
Causal version
We convert the officially available DeBERTa checkpoint into a HuggingFace (Wolf et al., 2020) implementation of AutoModelForCausalLM (following Section 2.1), and release it openly at https://huggingface.co/ltg/deberta-xxlarge-fixed.
Figure 4: The performance improvement with increased number of in-context examples. We compare the in-context learning ability of 1.4B DeBERTa (in red) with 1.3B GPT-3 (in blue) using prompts without any completed examples (0-shot), prompts with a single randomly sampled gold sample (1-shot), and prompts with ‘few’ examples (4 – 64 examples, depending on the task). This figure demonstrates that a masked language model behaves similarly to a causal language model in the in-context learning regime.
@@ 4 Evaluation
As our goal is to compare two language models released around the same time in 2020 – GPT-3 and DeBERTa – we replicate the evaluation setup used for GPT-3 (Brown et al., 2020) and apply it to the latter model. This also means that we follow GPT-3 and divide the tasks into generative ones (such as machine translation) and into classification tasks (such as BoolQ) – the first group uses the method described in Section 2.1 and the second type of task uses the ranking described in Section 2.2. Generation is performed with beam search (4 candidate beams), and ranking uses the modified PLL scores (and the normalized unconditional probability of completions $\frac{\operatorname{\mathbb{P}}(\text{completion}\ |\ \text{context})}{\operatorname{\mathbb{P}}({\text{completion}}\ |\ \text{answer context})}$ for ARC and OpenBookQA, again replicating the choices for GPT-3). We also use the exact same prompt templates, with the exception of the machine translation task – its template did not produce any meaningful output, and so we decided to use the simple prompt template from Garcia et al. (2023) instead. More details on the evaluation setup can be found in Appendix A. Note that using prompts optimized for GPT-3 is slightly unfair to all other models, as prompting has a strong influence on performance (Gonen et al., 2023), but we believe that it makes the results more convincing than if we were to do extensive prompt engineering.
@@
Before looking into the details of each group, we show the overall aggregated scores for each group in Figure 1 and Figure 4. The first figure shows how the performance of both models scales with their size, while the latter figure compares the in-context learning abilities of the two language models. We also provide a qualitative evaluation of text generation in Appendix B and full results in Appendix D.
Table 1: Natural language understanding results. All results in this table are evaluated with accuracy (higher is better). The table shows the performance of the largest available DeBERTa (1.4 billion parameters) and of a similarly-sized GPT-3 model, the best results are boldfaced. The average score is calculated over averaged task scores (in case a task uses more than one metric).

### v2

@@ Abstract
While in-context learning is commonly associated with causal language models, such as GPT, we demonstrate that this capability also ‘emerges’ in masked language models. Through an embarrassingly simple inference technique, we enable an existing masked model, DeBERTa, to perform generative tasks without additional training or architectural changes. Our evaluation reveals that the masked and causal language models behave very differently, as they clearly outperform each other on different categories of tasks. These complementary strengths suggest that the field’s focus on causal models for in-context learning may be limiting – both architectures can develop these capabilities, but with distinct advantages; pointing toward promising hybrid approaches that combine the strengths of both objectives.
@@ 1 Introduction
Masked language models used to dominate the field of natural language processing due to their adaptability across diverse tasks and their superior performance compared to causal language models (Radford et al., 2018; Devlin et al., 2019). Between 2018 and 2020, the field witnessed a surge in the development of these models (Devlin et al., 2019; Liu et al., 2019; Lan et al., 2020, inter alia). However, the field dramatically shifted with GPT-3 and its introduction of in-context learning – the ability to infer and perform tasks from prompts and examples without any finetuning (Brown et al., 2020). This capability eliminated the need for task-specific training data and deep-learning expertise, making such models far more practical for real-world applications. This perceived advantage led many researchers and practitioners to abandon masked language models in favor of GPT-style architectures.
@@
Figure 1: The average 1-shot performance across four groups of NLP tasks   We compare the scaling abilities of DeBERTa (four sizes in red) with GPT-3 (eight sizes in blue). Even though these models rely on different training objectives, they scale in a similar log-linear manner overall. Yet, on a task-by-task basis, the pretraining methods lead to substantial differences between them.
@@
Previous studies of ‘emergent’ in-context learning abilities have focused almost exclusively on causal language models, creating a widespread assumption that this capability is unique to them (Saunshi et al., 2021; Olsson et al., 2022; Wei et al., 2022; Wang et al., 2023, inter alia). In this paper, we challenge this assumption by demonstrating that in-context learning can emerge in masked language models as well. In-context learning is a more general phenomenon and should not be studied with a singular pretraining objective in mind. Moreover, the assumed inability of masked language models to perform (generative) in-context learning has rendered them outdated – as explicitly noted by Tay et al. (2023): “BERT-style models are very restricted in their generative capabilities. Because of the cumbersomeness of task-specific classification heads, we strongly do not recommend using this class of autoencoding models moving forward and consider them somewhat deprecated.”
@@
In this paper, we challenge these prevailing assumptions about masked language models (MLMs). We present empirical evidence showing that DeBERTa, an MLM released just one month after GPT-3, is equally adept at in-context learning. Our findings suggest that the capacity for in-context learning is not tied to the training objective, but can be achieved across different types of language models. To our surprise, we found that DeBERTa does not simply mimic the performance of GPT-3 – the two model behave very differently – DeBERTa is clearly much better on tasks such as language understanding, and, on the other hand, much worse on tasks such as closed-book question answering. This suggests that masked and causal language modeling are two complementary training objectives and that there is a great potential for a training method that combines the strengths of both objectives. Finally, scaling (performance improvement with increased size of pretrained language models) is a crucial feature of modern language models; we demonstrate that MLMs do scale on in-context learning (Figure 1).
@@
We introduce a simple inference technique that transforms an MLM into a generative model without any further training. Using publicly available DeBERTa checkpoints, we show that the MLM training objective not only provides a versatile way of encoding text, but is also competitive in text generation and text completion ranking. This claim is tested by following the same evaluation suite as GPT-3, speculating on an ‘alternative reality’ in which a masked language model is the first model reported to achieve the so-called ‘emergent’ in-context learning abilities. While other masked language models could potentially demonstrate similar capabilities, we deliberately target DeBERTa because of its large size and its length-generalization abilities. Ultimately, our goal is to demonstrate that MLMs can perform in-context learning and that they can be surprisingly good at doing so.
Outline
@@
Figure 2: Illustration of the proposed methods for using a masked language model for text generation and text ranking   We show how to adapt a masked language model for in-context-learning tasks through simple input reformatting, requiring no additional training. Left: Text generation is achieved by 1) appending [MASK] tokens to the input prompt, 2) predicting the next token for the first mask, and 3) iteratively appending new masks and predicting tokens. Right: A similar approach is used to retrieve a pseudo-log-likelihood score of a text sequence that can be used to rank multiple sequences by their individual likelihoods. Both methods maintain the model’s original architecture while enabling new capabilities through careful input formatting.
@@ 2.1 Text generation
This straightforward inference scheme would be enough if the pretraining process were designed with this use case in mind. However, since our goal is to repurpose an existing masked language model, we have to complicate the method with two modifications that are also illustrated in Figure 2:
@@
    However, the addition of this end-of-sequence token creates a problem – it raises the probability that the masked token should end the sequence (for example with a full stop). Thus, in order to obtain a less restricted continuation, we include additional [MASK] tokens to pad the space in front of the end-of-sequence token. Specifically, we use two additional masks for the DeBERTa models. [^1] Note that this is not an arbitrary number but it is model-specific – DeBERTa models were pretrained to unmask spans of masked tokens where the longest allowed spans are three tokens long (He et al., 2021). This decision is later ablated in Appendix B.
@@
While this method works with the same quadratic time complexity (in sequence length), it is slower in practice because it is not possible to cache the intermediate self-attention key and value vectors. Instead, these have to be recomputed every step due to the bidirectional nature of the model. While our current implementation prioritizes demonstrating the core capability over optimization, several promising approaches could address these computational limitations in future work. For example, using prefix language modeling or selectively updating hidden vectors could significantly improve efficiency. We leave these optimizations for future work to maintain focus on establishing the fundamental ability of MLMs to generate text.
@@ 2.2 Ranking
While this equation matches the training objective of causal language models, it is not suitable for masked language models because they are not trained to estimate $\operatorname{\mathbb{P}}(w_{i}\,|\,c\oplus w_{0}\dots w_{i-1})$. Instead, Wang and Cho (2019) proposed to modify Equation 1 to make it more appropriate for BERT-like models. Salazar et al. (2020) then empirically showed that the resulting pseudo-log-likelihood (PLL) score can be used to accurately rank text sequences by their likelihood. Specifically, the PLL score is approximately proportional to the conditional probability of a text sequence and is computed as:
@@
However, this approximation gets very inaccurate when there are strong local dependencies between tokens. As a counterexample, the estimated likelihood of the multi-token word ‘supercalifragilisticexpialidocious’ is seven orders of magnitude higher than that of the single-token word ‘super’, which is clearly an incorrect estimation of the relative frequencies of these words. [^2] This is because the tokenizer splits the long word into 9 subwords and each of them is assigned an almost certain likelihood given the bidirectional context. The largest 1.5B DeBERTa estimates the pseudo log-likelihood of the first word to $-2.1$ while the second (very common) word has pseudo log-likelihood of $-9.6$.
@@ 2.3 Length generalization
A potentially limiting factor of using BERT-like models is that they are typically pretrained on shorter sequences than causal language models – arguably because the training of modern causal models is already optimized for in-context learning, which requires processing of long few-shot prompts. DeBERTa is not an exception to such pretraining; it was only trained with a relatively short maximum sequence length of 512 tokens (He et al., 2021). Fortunately, the architecture of DeBERTa can easily process much longer sequences than seen during training due to its use of relative positional embeddings with logarithmic buckets (Raffel et al., 2020).
Figure 3: Length generalization measured with a ‘needle in a haystack’ benchmark    The $x$-axis indicates the total size of the ‘haystack’ and the $y$-axis indicates the position of the ‘needle’; the values show the average exact-match accuracy for a particular configuration. Unfortunately, GPT-3 is a closed-source model and the original version is not accessible, so we use an open-source replication of GPT-3, OPT by Zhang et al. (2022), which should perform similarly on this task because of the the same transformer architecture as GPT-3. In particular, it uses absolute positional encoding, which strictly limits any model from generalizing to longer inputs than trained on.
We measure the extent to which DeBERTa generalizes to longer sequences with the ‘needle in a haystack’ test from RULER (Hsieh et al., 2024). Specifically, in our formulation of this task, a random 6-digit number (needle) is hidden in a long collection of essays (haystack). We then measure the exact-match accuracy of retrieving the hidden number given two variables: the total sequence length and the position of the needle in the haystack (more details about the evaluation setup are given in Section E.1).
The results in Figure 3 demonstrate that DeBERTa generalizes to sequences well beyond its training length, which is enabled by its relative positional encoding. For comparison, we also show results from OPT (Zhang et al., 2022), which uses absolute positional encoding like GPT-3. As expected from models using absolute positional encoding, performance drops sharply beyond the training length. This comparison highlights the importance of positional encoding choice for length generalization, independent of whether the model is masked or causal. In practice, this observation means that DeBERTa should be able to handle as many task demonstrations as models trained with longer sequences.
@@ 3 DeBERTa family of language models
This study uses the largest openly available English masked language model, DeBERTa with 1.5 billion parametrs, and its smaller configurations – 0.1B, 0.4B and 0.9B (He et al., 2021). DeBERTa is an improved version of a BERT language model (Devlin et al., 2019) that uses an advanced attention mechanism with relative positional embeddings – apart from being trained on a larger corpus and with a larger number of training steps.
@@
Compared to GPT-3 and modern large language models, DeBERTa was pretrained on a relatively small and clean text corpus – totalling 78GB of data after deduplication, the corpus is comprised of the English Wikipedia (12GB), BookCorpus (6GB; Zhu et al., 2015), OpenWebText (38GB; Gokaslan and Cohen, 2019), and STORIES (31GB; Trinh and Le, 2019). This is almost an order of magnitude less data than what was used to pretrain GPT-3. Notably, our strong results – despite this data disparity – could suggest that masked language models are more data-efficient than causal models for developing in-context learning capabilities. This claim would however need to be evaluated with a comprehensive study. In comparison, GPT-3 uses 570GB of filtered CommonCrawl, WebText2 (roughly 26GB), two web-scraped book corpora (roughly 17GB and 76GB), and the English Wikipedia (roughly 4GB, estimated from Brown et al. (2020)).
@@
Interestingly, even though DeBERTa uses a substantially smaller training corpus, it is trained on more than three times more tokens than GPT-3 (1 trillion compared to 300 billion). [^3] This means that DeBERTa was trained on dozens of repetitions of its training corpus, not unlike other popular masked language models (Devlin et al., 2019; Liu et al., 2019) – suggesting that this type of models operates under different ‘training laws’ than causal language models (Muennighoff et al., 2023). However, the loss is computed only on 15% of tokens (150 billion) and it is not clear what would be the effective number of tokens used for pretraining. Nevertheless, the total compute used for training depends on the number of input tokens and it is roughly $8.0\cdot 10^{21}$ FLOPs for the 1.5B DeBERTa, and $2.4\cdot 10^{21}$ FLOPs for the 1.3B GPT-3.
Causal conversion for HuggingFace
We have converted the officially available DeBERTa checkpoint into a HuggingFace (Wolf et al., 2020) implementation of AutoModelForCausalLM (following the method in Section 2.1), and released it openly at https://hf.co/ltg/deberta-xxlarge-fixed. The weights of this model are exactly the same as the official release from microsoft/deberta-v2-xxlarge, but we have fixed some bugs found in the original modeling script in addition to implementing the text generation abilities. [^4] Specifically: 1) incorrect name of the output embedding weights in the checkpoint file, 2) non-functioning implementation of the enhanced mask decoder (EMD), and 3) missing truncation of the relative positional indices. Similarly, we have also converted the smaller DeBERTa models and released them as ltg/deberta-base-fixed, ltg/deberta-large-fixed, and ltg/deberta-xlarge-fixed.
@@ 4 Evaluation
As our goal is to compare two language models released around the same time in 2020 – GPT-3 and DeBERTa – we replicate the evaluation setup used for GPT-3 (Brown et al., 2020) and apply it to the latter model. This also means that we follow GPT-3 and divide the tasks into generative ones (such as machine translation) and into classification tasks (such as BoolQ) – the first group uses the method described in Section 2.1 and the second type of task uses the ranking described in Section 2.2. Generation is performed with beam search (4 candidate beams), and ranking uses the modified PLL scores (and the normalized unconditional probability of completions $\frac{\operatorname{\mathbb{P}}(\text{completion}\ |\ \text{context})}{\operatorname{\mathbb{P}}({\text{completion}}\ |\ \text{answer context})}$ for ARC and OpenBookQA), again replicating the choices for GPT-3). We also use the exact same prompt templates, with the exception of the machine translation task – its template did not produce any meaningful output, and so we decided to use the simple prompt template from Garcia et al. (2023) instead. More details on the evaluation setup can be found in Appendix E. Note that using prompts optimized for GPT-3 is slightly unfair to all other models, as prompting has a strong influence on performance (Gonen et al., 2023), but we believe that it makes the results more convincing than if we were to do extensive prompt engineering.
@@
Before looking into the details of each group, we show the overall aggregated scores for each group in Figure 1 and Figure 4. The first figure shows how the performance of both models scales with their size, while the latter figure compares the in-context learning abilities of the two language models. We also provide a qualitative evaluation of text generation in Appendix A and full results in Appendix F.
Figure 4: The performance improvement with increased number of in-context examples   We compare the in-context learning ability of 1.5B DeBERTa (in red) with 1.3B GPT-3 (in blue) using prompts without any completed examples (0-shot), prompts with a single randomly sampled gold sample (1-shot), and prompts with few examples (4 – 64 examples, depending on the task). This figure demonstrates that a masked language model behaves similarly to a causal language model in the in-context learning regime. More detailed few-shot evaluation is in Figure 5.
4.1 Language understanding (SuperGLUE)
[…]
Results
We show the resulting scores of evaluation with the same prompts as GPT-3 in Table 1 and Appendix D. DeBERTa clearly outperforms its contemporary and scales much more favorably than the family of GPT models (Figure 1). Interestingly, the average performance of the 1.5B DeBERTa gets close to the reported performance of the largest 175B GPT-3 ($68.4$ vs. $68.9$, 1-shot). However, this average score is still far from the performance of a finetuned DeBERTa, which is more than 20 percentage points higher (He et al., 2021); the average few-shot performance of DeBERTa is slightly better than a finetuned BERT-large (Devlin et al., 2019; Wang et al., 2019).
Table 1: Natural language understanding results   All results in this table are evaluated with accuracy (higher is better). The table shows the performance of the largest available DeBERTa (1.4 billion parameters) and of a similarly-sized GPT-3 model, the best results are boldfaced. The average score is calculated over averaged task scores (in case a task uses more than one metric).